# Open Agent Specification (Agent Spec) Technical Report

### v3

@@ 2.2 Components
In Agent Spec, components are the building blocks which cover the commonly used elements in agentic systems, such as (but not limited to) LLMs and tools.
@@
  • Type safety + Usage hints for GUIs
    • Example: if a component’s property is expecting an LLM, only allow LLMs to be connected to it.
    • Example: highlight valid connections from one component to others in a graphical programming interface for agents.
  • Static analysis + validation
    • Examples: if a component is a Flow, it must have a start node. If a component is an Agent, it must have an LLM.
  • Ease of programmatic use
    • Component families each have corresponding class definitions in the Agent Spec SDK. This allows consumers (agent execution environments, editing GUIs) to utilize concrete classes, rather than having to implicitly understand which type has which properties.
@@ 2.2.1 Base Component
The basic block in Agent Spec is a Component, which by itself can be used to describe any instance of any type, guaranteeing flexibility to adapt to future changes.
Note that Agent Spec does not need to encapsulate the implementation code it describes; it simply needs to be able to express enough information to instantiate a uniquely identifiable object of a specific type with a specific set of property values.
A metadata field contains all additional information that could be useful in different usages of the Agent Specification. For example, GUIs will require including some information about the components (e.g., position coordinates, size, color, notes, …), so that they will be able to visualize them properly even after an export-import process. Metadata is optional and set to null by default.
@@
This type of relationship is applicable to both Agent components as well as Flow components (a fixed flow component may also use an LLM or memory component).
@@
Components might need some input data to perform their task and expose some output as result. These inputs and outputs must be declared and described in Agent Spec, so that users, developers and other components of the agent are aware of what is exposed as inputs and outputs respectively. Note that we do not add input/output schemas directly to the base Component class, as there are a few cases where they do not really apply (for example LLMConfig and edges, please see the next sections).
@@
Input and output schemas are used to define which values the different components accept as input or provide as output. We call these values "properties".
The term "properties" originates from JSON schema [8]. In the Agent Spec specification, we rely on this widely accepted and consolidated format. For more information about JSON schema definition, please check the official website at https://json-schema.org/understanding-json-schema.
In Agent Spec, a valid schema requires specific attributes:
  • title: the name of the property
  • type: the type of the property, following the type system described below
Additionally, users can specify:
  • default: the default value of this property
  • description: a textual description for the property
These are the minimal set of attributes (or annotations, as per JSON schema terminology) that we require to be supported for Agent Spec compatibility.
Type system
We rely on the typing system defined by the JSON schema format (see https://json-schema.org/understanding-json-schema/reference/type). At its core, JSON Schema defines the basic types available in most programming languages, though they may go by different names.
@@
In case of nested components (for example using an Agent component inside a Flow, or a Flow inside another Flow, or just a step in a Flow), the wrapping component is supposed to expose a (sub)set of the inputs/outputs provided by the inner components together with additional inputs/outputs it might generate as shown in Fig. 2. We require to replicate the JSON schema of each value in the inputs/outputs lists of every component that exposes it, hence, if a wrapper component exposes some inputs and outputs of some of its internal components, it will have to include them in its own list of inputs/outputs as well. This makes the specification of the components a bit more verbose, but clearer and more readable. Parsers of the specification are required to ensure the consistency of the input/output schemas defined in it.
Figure 2: Inputs and outputs of nested components, as well as those of their wrapping component
In the example above the wrapping Flow component (see sections below for more details about Flows) requires two inputs that coincide with two of the inputs exposed by the nested components and exposes one output among those exposed by the inner components in addition to one it computes itself.
@@
Some input and output schemas are automatically generated by components based on their configuration. However, for the sake of clarity and readability of Agent Spec, we require that the descriptions of inputs and outputs are always reported in the Agent Specification. Consequently, when a serialized specification is read and imported by a runtime or an SDK, the inputs and outputs configuration of each component must be validated against the one generated by its configuration (if any).
@@
Some components might infer inputs from special parts of their attributes. For example, some nodes (e.g., the LLMNode) extract inputs from their configuration attributes, so that their values can be adapted based on previous computations We let users define placeholders in some attributes, by simply wrapping the name of the property among a double set of curly brackets. The node automatically infers the inputs by interpreting those configurations and the actual value of the input/output will be replaced at execution time. Whether an attribute accepts placeholders depends on the definition of the component itself. For example, setting the prompt_template configuration of an LLMNode to "You are a great assistant. Please help the user with this request: {{request}}" will generate an input called "request" of type string.
@@ 2.2.2 Agent
The Agent is the top-level component that holds shared resources such as conversation memory and tools. It also represents the entry point for interactions with the agentic system. Its purpose is to fill out the values for all the properties defined in the outputs attribute, with or without any interaction with the user.
It is important to have a separate definition for Agents as components, so that the same Agent component can be reused several times in a Flow or as part of a more complex agent design, without replicating its definition multiple times.
@@ 2.2.3 LLM
Large Language Models (LLM) are used in agents as generative models. To embed them into the agent, they need to be configurable flexibly with configuration parameters like the connection details, model id, the generation parameters, and more.
By default, only the generation parameters used when prompting the LLM must be specified. Specific extensions of LLMConfig for the most common model providers can be derived.
Indeed, this is not limited to locally hosted models; in principle, it also enables the configuration of API endpoints such as those provided by OpenAI, Anthropic, Mistral, and others.
@@ 2.2.4 Tool
A tool is a procedural function or a Flow that can be made available to an Agent component to execute and perform tasks. The Agent component can decide to call a tool based on its signature and description. In a fixed Flow context, a node would need to be configured to call a specific tool.
In Agent Spec, we differentiate tools according to where the actual functionality is executed or run and specify three corresponding types:
  • ServerTools are executed in the same runtime environment where the agent is being executed. The definitions of these tools therefore must be available to the agent’s environment. It is expected that this type of tool will be limited in number, and generally in functionality.
  • ClientTools are not executed by the runtime. The client environment must execute the tool and provide the results back to the runtime (like OpenAI’s function calling model).
  • RemoteTools are run in an external environment and are triggered by an RPC or REST call from the agent runtime.
Agent Spec is not supposed to provide an implementation of how these types of tools should work but provides their representation, so that they can be correctly interpreted and ported across platforms and languages. In principle a Tool is a function that is called providing some parameters (i.e., inputs), performs some transformation, and returns the result (i.e., outputs). The function has a name, and a description that can be used by an LLM for context when it is relevant to perform a task. Therefore, we define Tool as an extension of ComponentWithIO.
Tools may specify multiple outputs. In this case, the expected return value of the tool is a dictionary where each entry corresponds to an element specified in the output. The key of the dictionary entry must match the name of the property specified in the output. The runtime should parse the dictionary to extract the different outputs and bind them correctly. While ServerTool and ClientTool do not require specific additional parameters, the
RemoteTool requires including also the details needed to invoke the remote tool call.
An important security consideration for Agent Spec is the exclusion of arbitrary code from the agent specification. The specification is limited to a comprehensive description of the tool—including its attributes, inputs, outputs, and other relevant details—but does not contain any executable code.
@@ 2.2.5 Flow
Flows (or graphs) are directed, potentially cyclic workflows. They can be thought of as "subroutines" that encapsulate consistently repeatable processes. As such, they provide more determinism and reliability compared to puristic Agent components at the cost of flexibility.
Flows are executable as tools by Agent components. Each Flow requires 0 or more inputs (any input edges expressed on the starting node of the graph) and may produce 0 or more outputs (any output edges expressed by the terminal node of the graph - note that a graph can have more than one terminal node, in the case of branching logic). As such, the I/O surface of a Flow is directly mutually exchangeable with that of a tool.
Flow objects specify the the directed graph’s entry point, nodes, and edges. Components inherent in the Flow are described in subsequent sections.
A Flow exposes the set of outputs that are available in all EndNode objects, and all inputs that are defined in its StartNode.
@@
For Flows, Agent Spec defines separate relationships (edges) for both properties and control flow. This allows Agent Spec to unambiguously handle patterns such as conditional branches and loops, which are challenging to support with data-only flow control.
Relations are only applicable for Flows. All relations express a transition from some source node to some target node.
@@ 2.2.6 ControlFlowEdge
A control relationship defines a potential transition from one branch of a node to another. The actual transition to be taken during a given execution is defined by the implementation of a specific node.
The from_branch attribute of a ControlFlowEdge is set to null by default, which means that it will connect the default branch (i.e., “next” ) of the source node.
@@ 2.2.7 DataFlowEdge
Considering an I/O system, a component id alone is not sufficient to identify a data relationship. To this purpose, a data relationship is needed to define what output property value should be used to fill an input property.
A component may take multiple input parameters and/or produce multiple output values. We must be able to determine what value or reference is mapped to each input and where each output is used. Destinations (Inputs of the destination node) can be connected to outputs of another node or to static values. Sources (Outputs of the source node) can be connected to inputs of another node or left unconnected.
@@
(a) Control flow with two branches that converge into the same node with a property filled based on the branch taken
@@
Figure 3: Possible connections formed by ControlFlowEdges and DataFlowEdges. ControlFlowEdges are shown as solid lines, DataFlowEdges with dotted lines.
In the definition of a Flow, it is allowed to connect two or more different data outputs to the same input as shown in Fig. 3. The most recently executed node with an output connected to the input takes precedence. In other words, the behavior is like having the input exposed as a public variable and every node that has an output connected to that input updates its value.
For the control flow, instead, edges just define which are the allowed "next step" transitions: multiple connections from the same outgoing control flow branch are not allowed, and they do not involve any ambiguity (note that we do not enable parallelism through edges).
Optionality of data relationships
Some frameworks do not require specifying the data flow explicitly as they assume that all the properties are publicly exposed in a shared "variable space" that any component can access, and the read/write strategy is simply name based. In practice, this means that:
  • When a component has an input with a specific name, it will access the public variable space to read the value of the variable with that name.
  • When a component has an output with a specific name, it will write (or overwrite if it already exists) into the public variable space the variable with that name.
We let users adopt this type of I/O system by setting the data_flow_connections parameter of the Flow to None. In this case, the name-based approach explained above will be adopted, by using the shared variable space described above as the public space where values are published.
Note that this name-based approach can be expressed by explicitly defining data-flow connections (but the opposite is not possible) by connecting all the outputs to the inputs with the same name, following the control flow connections in the Flow to account for overwrites.
Flow state
We define as "flow state" the set of all values that describe the current stage of execution within the flow (the point of execution reached). Users have the ability to specify which set of values within the flow state are publicly exposed to all nodes.
These values are treated differently compared to values that are exposed through inputs and outputs of nodes:
  • They are always available in every node and do not need to be connected by a DataFlowEdge.
  • If a component writes an output which name matches a flow state value, the respective value in the flow state is overwritten.
Conversation
At the core of the execution of a conversational agent, there is the conversation itself. The conversation is implicitly passed to all the components throughout the flow. It contains the messages being produced by the different nodes: each node in a flow and each agent can append messages to the conversation if needed. Messages have, at least, a type (e.g., system, agent, user), content, the sender, and the recipient.
@@ 2.2.8 Node
A Node is a vertex in a Flow and is derived from ComponentWithIO. Agent Spec provides a library of nodes designed to streamline workflow orchestration, enabling rapid development and seamless integration for a wide range of applications:
  • LLMNode: uses a LLM to generate some text from a prompt.
  • APINode: performs an API call with the given configuration and inputs.
  • AgentNode: runs a multi-round conversation with an Agent component, used to better structure agents and easily reuse Agent components.
  • FlowNode: runs a Flow, used to better structure agents and easily reuse Flows.
  • MapNode: performs a map-reduce operation on a given input collection, applying a specified Flow to each element in the collection.
  • StartNode: entry point of a Flow.
  • EndNode: exit point of a Flow.
  • BranchingNode: allows conditional branching based on the value of an input.
  • ToolNode: executes a tool.
A more detailed description of each node type is given in https://oracle.github.io/agent-spec/agentspec/language_spec_25_4_1.html#standard-library-of-nodes-to-use-in-flows.
@@ 2.3 Open Agent Specification Model
The objective of Agent Spec is to be agnostic with respect to any agentic framework and the programming language used to build and execute an agent. For this reason, an agent represented with Agent Spec should be exportable to, and importable from a representation that is specified in a common serialization language. To this end, Agent Spec adopts JSON as the designated language for the serialization of its components.
All building blocks of Agent Spec are designed with the goal of being trivially serializable. It follows that the Agent Specification of a component can be obtained by serializing its definition according to the JSON specification. In particular, the serialization of a component includes its attributes, plus an additional field called component_type that will be used at deserialization time to infer the correct component type to deserialize.
Note that, as defined in the Agent Spec language, components can be declared only once in the representation and then referenced according to the rules described in the Symbolic references section of this document.
A few examples of component JSON serializations are given at https://oracle.github.io/agent-spec/agentspec/language_spec_25_4_1.html#language-examples.

### v4

@@ 4.2 Components
In Agent Spec, components are the building blocks which cover the commonly used elements in agentic systems, such as (but not limited to) Agents, LLMs, tools, etc.
@@
  • Type safety and GUI guidance: For example, if a component property expects an LlmConfig, only compatible extensions can be connected, or valid connections can be highlighted to improve the agent-building experience in graphical interfaces.
  • Static analysis and validation: For instance, if a component is a Flow, it must include a start node; if it is an Agent, it must contain an LLM.
  • Ease of programmatic use: Component families have corresponding class definitions in the Agent Spec SDKs (see Section 5). This allows consumers (e.g., execution environments or GUI editors) to work with concrete classes directly, rather than inferring which properties belong to which types.
@@ 4.2.1 Base Component
The basic block in Agent Spec is a Component, which by itself can be used to describe any instance of any type, guaranteeing flexibility to adapt to future changes. It is important to note that Agent Spec does not need to encapsulate the implementation code it describes; it simply needs to be able to express enough information to instantiate a uniquely identifiable object of a specific type with a specific set of property values.
A metadata field contains all additional information that could be useful in different usages of the Agent Specification. For example, GUIs will require including some information about the components (e.g., position coordinates, size, color, notes), so that they will be able to visualize them properly even after an export-import process. Metadata is optional and set to null by default.
@@
"\$component_ref:{COMPONENT_ID}"
This type of relationship is applicable to both Agent components as well as Flow components (a workflow component may also use an LLM or memory component).
@@
Components might require input data to perform their task and produce output as result. These inputs and outputs must be declared and described in Agent Spec, so that users, developers and other components of the agent are aware of what is exposed as inputs and outputs respectively. Note that input/output schemas are not added directly to the base Component class, as there are a few cases where they do not apply (for example LLMConfig and edges; see the following sections for details).
@@
Input and output schemas define the "values" that different components can accept as input or produce as output. These values are referred to as "properties" based on JSON schema standard [jsonschema].
@@
In the case of nested components (for example, using an Agent component inside a Flow, or a Flow inside another Flow, or just a step in a Flow), the wrapping component is supposed to expose a (sub)set of the inputs/outputs provided by the inner components together with additional inputs/outputs it might generate as shown in Fig. 3.
Figure 3: Illustration of input/output exposure in nested components.
Each value listed in the inputs and outputs must replicate its corresponding JSON schema in every component that exposes it. Consequently, if a wrapper component makes available certain inputs or outputs from its internal components, these must also be explicitly included in the wrapper’s own input/output lists. This makes the specification of the components more verbose, but transparent and more readable. Parsers of the specification are required to ensure the consistency of the input/output schemas defined in it.
In the example above, the wrapping Flow component (see sections below for more details on Flows) requires two inputs that match those exposed by its nested components and produces one output derived from the inner components, along with an additional output computed internally.
@@
Some input and output schemas are automatically generated by components based on their configuration. However, for the sake of clarity and readability of Agent Spec, all input and output descriptions must be explicitly included in the Agent Specification. Consequently, when a serialized specification is loaded by a runtime or SDK, the inputs and outputs configuration of each component must be validated against those automatically generated from its configuration (if applicable).
@@
Some components can infer inputs from specific parts of their configurations. For example, certain nodes (e.g., the LLMNode) extract inputs from their configuration attributes, allowing their values to be dynamically adapted based on previous computations. Users can define placeholders within these attributes by enclosing the name of a property in double curly brackets. The node automatically infers the inputs by interpreting those configurations and the actual value of the input/output will be substituted at execution time. Whether an attribute accepts placeholders depends on the definition of the component itself. For example, setting the prompt_template configuration of an LLMNode to "You are a great assistant. Please help the user with this request: {{request}}" will generate an input called "request" of type string.
@@ 4.2.2 Agent
The Agent is the top-level component that holds shared resources such as conversation memory and tools. It also serves as the entry point for interactions with the agentic system. Its primary goal is to populate the values of all properties defined in its output attribute, either through user interaction or autonomous computation.
Defining Agents as standalone components is essential, as it enables the same Agent component to be reused across multiple Flows or within more complex agent architectures, without the need to duplicate its definition.
@@ 4.2.3 LLM
Large Language Models (LLMs) are used within agents as generative components. To integrate them effectively, they must be flexibly configurable through parameters such as connection details, model identifiers, generation settings, and other options.
By default, only the generation parameters required for prompting the LLM need to be specified. Specific extensions of LLMConfig can be derived to support the most common model providers. This approach is not limited to locally hosted models; it also supports the configuration of external API endpoints, such as those offered by OpenAI, Anthropic, Mistral, and others.
@@ 4.2.4 Tool
A tool is a procedural function or a Flow that can be made available to an Agent component to execute and perform tasks. The Agent can decide which tool to invoke based on the tool’s signature and description. In a Flow context, a node must be explicitly configured to call a specific tool.
A novelty in Agent Spec is the classification of tools according to where their actual functionality is executed. Four corresponding types are defined:
  • ServerTools are executed within the same runtime environment as the agent itself. The definitions of these tools therefore must be available in the agent’s environment. Such tools are typically limited in number and functionality.
  • ClientTools are not executed by the runtime. Instead, they run in the client environment, which is responsible for executing the tool and returning the results to the runtime (similar to OpenAI’s function calling model).
  • RemoteTools are executed in an external environment and triggered through an RPC or REST call initiated by the agent runtime.
  • MCPTools enable execution of tools provided by a MCP server in accordance with the MCP specification [mcp].
Agent Spec does not prescribe how these tool types should be implemented. Rather, it provides a representation format that allows them to be correctly interpreted and exchanged across different platforms and programming languages.
Conceptually, a Tool is a function that accepts parameters (inputs), performs a transformation, and returns a result (outputs). Each tool has a name and a description, which can provide context to an LLM when deciding which tool to use for a given task. Consequently, a Tool is defined as an extension of ComponentWithIO.
Tools may produce multiple outputs. In such cases, the expected return value is a dictionary, where each key corresponds to the name of an output property defined in the tool’s schema. The runtime should parse the dictionary to extract the different outputs and bind them correctly.
While ServerTool and ClientTool do not require additional parameters beyond their standard definition, a
RemoteTool must include the necessary details for invoking the remote call.
Finally, an important security consideration in Agent Spec is the exclusion of arbitrary code from the specification. The Agent Specification only provides a structured description of each tool, including its attributes, inputs, outputs, and metadata but does not contain any executable code.
@@ 4.2.5 Flow
Flows (or graphs) are directed, potentially cyclic workflows. They can be regarded as "subroutines" that encapsulate structured sequences of operations. As such, Flows offer greater determinism and reliability compared to purely agent-driven components at the cost of reduced flexibility.
Each Flow may require zero or more inputs (corresponding to the inputs expressed on the starting node of the graph) and may produce zero or more outputs (represented by the terminal node of the graph - note that a graph might contain multiple terminal nodes in cases involving branching logic).
Flow objects specify the directed graph’s entry point, its nodes, and the edges connecting them. The individual components that compose a Flow are described in detail in the following sections.
A Flow exposes all inputs defined in its StartNode and all outputs that are available across its EndNode objects.
Conversation
At the core of a conversational agent’s execution lies the conversation itself. The conversation is implicitly propagated across all components within the Flow and maintains the collection of messages generated throughout execution. Each node in a Flow, as well as each Agent, might append messages to this conversation as needed. Every message contains, at minimum, the following attributes: role (e.g., system, agent, user), content, sender, and recipient.
Input/Output Data Space
Alongside the conversation, flows have another data space composed of inputs and outputs generated by nodes. When a node generates an output, this is exposed by the node itself, so that other nodes can consume it as input, if needed. The input-output connections, called data flow, can be defined through explicit data relationships, or through implicit, name-based property matching, as depicted in the following sections.
@@
Within Flows, Agent Spec distinguishes between two categories of relationships (or edges): those governing data flow and those governing control flow. This distinction enables Agent Spec to unambiguously represent complex execution patterns such as conditional branches and loops, which are difficult to model using data flow alone.
Relationships are defined exclusively within the context of Flows. Each relationship expresses a directed transition from a source node to a target node.
@@ 4.2.6 ControlFlowEdge
A control flow relationship specifies a potential transition from one branch of a node to another. The actual transition followed during execution is determined by the implementation of the corresponding node.
The attribute from_branch of a ControlFlowEdge is set to null by default, indicating that the edge connects to the node’s default branch (i.e., next).
@@ 4.2.7 DataFlowEdge
In an input/output system, a component identifier alone is insufficient to fully specify a data relationship. Therefore, Agent Spec defines explicit data flow relationships to describe how the output property of one node is mapped to the input property of another.
A component might accept multiple input parameters and/or produce multiple output values. Hence, it is necessary to define which output or reference is associated with each input, and conversely, how each output is consumed.
Destinations (the inputs of the target node) can be connected to outputs from another node or assigned to static values. Conversely, sources (the outputs of the source node) may be connected to inputs of subsequent nodes or left unconnected if unused.
@@
(a) Control flow with two branches that converge into the same node with a property filled based on the branch taken.
@@
Figure 4: Control- and data-flow relationships in Agent Spec. Solid lines denote control-flow transitions, while dotted lines indicate data propagation.
In the definition of a Flow, it is permissible to connect two or more distinct data outputs to the same input as shown in Fig. 4. In such cases, the value produced by the most recently executed node among those connected to the input takes precedence. Conceptually, this behavior is analogous to exposing the input as a public variable, where each connected node can update its value upon execution.
In contrast, control flow edges define the set of permitted next-step transitions. Multiple outgoing connections from the same control flow branch are not permissible, ensuring unambiguous execution order. It should also be noted that parallelism is not enabled through control flow edges.
Optionality of data flow relationships
Some frameworks do not require the explicit specification of data flow, as they assume that all properties are publicly exposed within a shared "variable space" accessible to any component. In such systems, read and write operations are determined purely by name-based resolution.
In practice, this means that:
  • When a component defines an input with a given name, it retrieves the value of the variable with the same name from the shared variable space.
  • When a component defines an output with a given name, it writes (or overwrites, if already present) the corresponding variable in the shared variable space.
Users can adopt this type of I/O mechanism in Agent Spec by setting the data_flow_connections parameter of a Flow to None. In this configuration, the system follows the name-based approach described above, using the shared variable space as the public medium for publishing and accessing values.
It is worth noting that this name-based strategy can be expressed equivalently by explicitly defining the corresponding data flow connections by connecting all outputs and inputs that share the same name and following the control flow to account for overwrites. However, the converse is not true: an explicitly defined data flow configuration cannot always be reduced to a purely name-based representation.
@@ 4.2.8 Node
A Node represents a vertex within a Flow and is derived from the ComponentWithIO class. Agent Spec provides a standard library of node types designed to facilitate workflow orchestration, enabling efficient development and seamless integration across a wide range of applications.
The main node types include:
  • LLMNode: Utilizes a large language model (LLM) to generate text from a given prompt.
  • APINode: Performs an API call based on the specified configuration and inputs.
  • AgentNode: Executes a multi-turn interaction with an Agent component, supporting modular agent design and reuse.
  • FlowNode: Executes a Flow, enabling hierarchical structuring and reuse of Flows within more complex architectures.
  • MapNode: Performs a map-reduce operation over an input collection, applying a specified Flow to each element in the collection.
  • StartNode: Defines the entry point of a Flow.
  • EndNode: Defines the termination point of a Flow.
  • BranchingNode: Enables conditional branching based on the value of a specified input.
  • ToolNode: Invokes and executes a defined Tool.
  • InputMessageNode: Interrupts temporarily the execution to retrieve user input.
  • OutputMessageNode: Appends an agent message to the conversation.
A comprehensive description of each node type is provided in the Agent Spec documentation, available at [language_spec_25_4_1].
@@ 4.3 Open Agent Specification Model
The primary objective of Agent Spec is to remain agnostic with respect to any agentic framework and the programming language used to build and execute an agent. Consequently, an agent represented in Agent Spec should be exportable to, and importable from a common, framework-independent serialization format. To achieve this, Agent Spec adopts JSON as the designated language for the serialization of its components.
All building blocks within Agent Spec are explicitly designed to be trivially serializable. Thus, the Agent Specification of any component can be directly obtained by serializing its definition according to the JSON standard. Each serialized component includes its set of attributes, along with an additional field, component_type, which is used during deserialization to infer the appropriate component type.
As defined by the Open Agent Specification, each component may be declared only once within a representation and subsequently referenced symbolically, following the rules outlined in the Symbolic References section of this article. Illustrative examples of component serializations in JSON format are available at [language_spec_25_4_1].